What is the value of z in the solution to this system of linear equations?
7y – 11z = 6
6y + 2z = 28

Answer by edjones(8007)

7y – 11z = 6
6y + 2z = 28
.
2z=28-6y
z=14-3y
.
7y-11(14-3y)=6
7y-154+33y=6
40y=160
y=4
.
z=14-3*4
=14-12
=2
.
Ed